random variable X has this probability distribution:


x 0 1 2 3 4 5


P(X=x) 0.1 0.2 0.05 0.1 ? 0.3


Find f(4)



1
Expert's answer
2022-04-21T03:23:49-0400

The sum of all probabilities for different values of "x" must be equal to "1." We receive: "\sum_{x=0}^5P(X=x)=1". From the latter we get: "0.1+0.2+0.05+0.1+0.3+P(X=4)=1". We receive: "0.75+P(X=4)=1;" "P(X=4)=0.25". In case we denote: "f(x)=P(X=x)", we get: "f(4)=P(X=4)=0.25."